OpenStudy (anonymous):

A solution of cough syrup contains 5.00\% active ingredient by volume. If the total volume of the bottle is 18.0 mL, how many milliliters of active ingredient are in the bottle?

OpenStudy (anonymous):

Once again, it's simple problem of percentage. If the whole thing is 18 mils and 5% of it is active ingredient, then all you have to do is multiply the total volume by .05 (which is 5%).

OpenStudy (anonymous):

So the answer is 9

OpenStudy (anonymous):

I need more help I still don't understand :(

OpenStudy (anonymous):

Nope it's .9mL

OpenStudy (anonymous):

Do you understand percentages? 50%=.5=1/2 5%=.05=1/20

OpenStudy (anonymous):

Sorry but I am really bad at percentage and word problems :(

OpenStudy (anonymous):

Whenever a percentage is given when it comes to a whole, multiply it by the percentage's decimal counterpart, which in this case is .05. An easy way to remember that is to remember that .05 is 5% of 1.

OpenStudy (anonymous):

18mL(.05)=.9mL

OpenStudy (anonymous):

I see want you did but I have a hard problem than that. A patient needs to take 625 mg of ibuprofen twice daily. The pills in the bottle are each 250 mg. How many pills does the patient need to take each time she wants the 625 mg dosage? Assume that these pills are scored and can be cut in half if necessary.

OpenStudy (anonymous):

So ow do you start the problem off

OpenStudy (anonymous):

Brandon are you still there I need your help

OpenStudy (anonymous):

I am here, just distracted :P

OpenStudy (anonymous):

ohh well do you understand this problem bc you are very helpful

OpenStudy (anonymous):

It's very simple. 625mg is the total dosage and each individual pill is 250mg. So that's two pills to get 500mg and half of another (125mg) to get 625mg. Your answer is 2.5.

OpenStudy (anonymous):

omg when you explain it makes sense

OpenStudy (anonymous):

bc 250+250+125=625 and that's 2 and a half a pill

OpenStudy (anonymous):

Medal? :D

OpenStudy (anonymous):

You deserve like 3 medals, 8P